These were accepted as a reasonable account of the last meeting.
See Inf report
See Inf report
Change WHO to ALL - comments/opinions to Toby before next Ops meeting.
Next pick up is Thu 17th November. Please do not put any kit into B.03 before then. Services Unit need to organise removal of old tape library from KB server room.
A reminder that these need to be made a priority.
Booked for Tuesday 13th December. £10 deposit needs to be given to Neil as soon as possible.
Gordon has written initial page but more to be added.
Waiting for new version of dsu
Neil has produced report so Alastair/CEG need to make decision.
Gordon will investigate after completing MacOS documentation.
Only a few left!
Stalled again for 6 months. Waiting on clarity on UPS's
See US report
Blog articles discussed:
Is this still required ?
Blog articles completed:
Report from Computing Executive Group
Reports from Units
Non-/24 files should set netmask_eth0 and perhaps also broadcast_eth0..... George has asked the question "Is there a better way to do this ? If you have any comments/suggestions please reply to George and MPU.
This should be a priority.
As soon as a new version of dsu comes out, older versions immediately stop functioning. However it's possible that the new version may support RHEL7.3 rather than 7.2 so advice is to get going sooner rather than later!
Note there is a big exam on Tuesday 20th December. We may need volunteers to help out on the day.
Topics for discussion
We had an open discussion, the main points coming from this being:
The next meeting will be on Wednesday 23rd November 2016 at 10:00 in IF-4.31, chaired by Alison.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh